Question 1: Let A and B be two events. Determine whether the below statements are true or false. Give a proof (if true) or a counterexample (if false).

a) P(A|B) + P(A|Bc) = 1

b) P(Ac|B) + P(A|B) = 1
Answer #1

a) False ; (correct statement is    P(A|B)+P(Ac |B) =1 )

b) true

since P(Ac |B)+P(A|B) =P(Ac n B)/P(B)+P(A n B)/P(B) =(1/P(B))*(P(A n B)+P(Ac n B)) =(1/P(B))*P(B) =1